# Field of values analysis that includes zero for preconditioned nonsymmetric saddle-point systems

### v1

@@ 4.2 Stokes-Darcy Equations
[…]
where $\mathbf{u}$ satisfies the incompressibility condition $\nabla\cdot\mathbf{u}=0$. Here, $\Omega_{S}$ and $\Omega_{d}$ are assumed to be simple domains, e.g., the unit squares in two dimensions, with a polygonal interface. The operator $D$ is defined as $D(\mathbf{u})=\frac{1}{2}(\nabla\mathbf{u}+\nabla\mathbf{u}^{T}).$ The physical parameters $\nu$ and $k$ denote the viscosity coefficient and hydraulic constant, respectively. The constant $G$ represents an experimentally-determined constant related to the Beavers-Joseph-Saffman interface condition. Finally, $\mathbf{n}_{12}$ and $\mathbf{\tau}_{12}$ are unit normal and tangential vectors; see [4] for details. We use the finite element discretization described in [4, 9, 5]; some details on the Stokes part are similar to Section 4.1. Full details on the discretization of the entire problem are omitted since this is not the focus of our paper. The discretization yields the following linear system:
@@
Then, assuming that $\nu$ is sufficiently large (which corresponds to requiring $\alpha$ to be sufficiently small in our analysis in Section 3; see (9)), let us define
@@
We use the following example from [4]. We choose $\Omega_{s}$ to be $[0,1]^{2}$ and $\Omega_{d}$ to be $[0,1]\times[1,2]$. $\Gamma_{n}$ is $\{x=0,y\in[0,1]\}\cup\{x=1,y\in[0,1]\}$. Boundary conditions and right-hand side are computed from the following exact solution:
@@
Table 2: Iteration counts for the Stokes-Darcy equations.
For the diagonal preconditioner, we have observed experimentally for the smaller-size problems that the parameters of Lemma 2.4 satisfy $b\approx 9.18$, $c\approx 0.08$, and $bc<1$. For the upper triangular preconditioner, $b\approx 8.28$, $c\approx 0.11$ and $bc<1$.
@@ 5 Concluding Remarks
When applying Theorem 1.5, a disk must be excluded from the field of values, and the remaining part should not surround the origin, as we have illustrated in Figure 1. To accomplish this, we require the imaginary part of the FOV to be small enough, which means that for the nonsymmetric saddle-point systems we consider, the preconditioned matrix is close to normal, or equivalently, the skew-symmetric part of preconditioned operator needs to be small in norm.
A finer geometric study of the field of values, beyond bounding it just by using the imaginary axis, may allow for loosening the aforementioned restriction. It would be useful to enhance the set of nonsymmetric saddle-point linear systems for which FOV analysis of the type we have offered is applicable.

### v2

@@ 4.2 Stokes-Darcy Equations
$-\nabla\cdot(2\nu D(\mathbf{u})-p\mathbf{I})$ $=\mathbf{f}^{s}$ $\text{ in }\partial\Omega_{s},$
$\nabla\cdot\mathbf{u}$ $=0$ $\text{ in }\partial\Omega_{s},$
$\mathbf{u}$ $=\mathbf{g}^{s}$ $\text{ on }\Gamma_{s}=\partial\Omega_{s}\cap\partial\Omega,$
$-k\Delta\phi$ $=f^{d}$ $\text{ in }\Omega_{d},$
$\phi$ $=g^{d}$ $\text{ on }\Gamma_{d},$
$k\nabla\phi\cdot\mathbf{n}$ $=g^{n}$ $\text{ on }\Gamma_{n},$
$\mathbf{u}\cdot\mathbf{n}_{12}$ $=-k\nabla\phi\cdot\mathbf{n}_{12}$ $\text{ on }\Gamma_{I},$
$(-2\nu D(\mathbf{u})\cdot\mathbf{n}_{12}+p\mathbf{n}_{12})\cdot\mathbf{n}_{12}$ $=\phi$ $\text{ on }\Gamma_{I},$
$\mathbf{u}\cdot\boldsymbol{\tau}_{12}$ $=-2\nu G(D(\mathbf{u})\mathbf{n}_{12})\cdot\boldsymbol{\tau}_{12}$ $\text{ on }\Gamma_{I},$
where $\mathbf{u}$ satisfies the incompressibility condition $\nabla\cdot\mathbf{u}=0$. Here, $\Omega_{S}$ and $\Omega_{d}$ are assumed to be simple domains, e.g., the unit squares in two dimensions, with a polygonal interface. The operator $D$ is defined as $D(\mathbf{u})=\frac{1}{2}(\nabla\mathbf{u}+\nabla\mathbf{u}^{T}).$ The physical parameters $\nu$ and $k$ denote the viscosity coefficient and hydraulic constant, respectively. The constant $G$ represents an experimentally-determined constant related to the Beavers-Joseph-Saffman interface condition. Finally, $\mathbf{n}_{12}$ and $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{\tau}_{12}$} are unit normal and tangential vectors; see [4] for details.
We use the finite element discretization described in [4, 9, 5]; some details on the Stokes part are similar to Section 4.1. We note that there are several other distinct possibilities here for different discretizations; see, e.g., [19]. Full details on the discretization of the entire problem are omitted since this is not the focus of our paper. The discretization yields the following linear system:
@@
Then, assuming that $\nu$ is sufficiently large (which corresponds to requiring $\eta$} to be sufficiently small in our analysis in Section 3; see (11)), let us define
@@
We use the following example from [4]. We choose $\Omega_{s}$ to be $0,1]^{2}$ and $\Omega_{d}$ to be $0,1]\times[1,2]$. $\Gamma_{n}$ is $\{x=0,y\in[0,1]\}\cup\{x=1,y\in[0,1]\}$. Boundary conditions and right-hand side are computed from the following exact solution:
@@
Table 2: Iteration counts for the Stokes-Darcy equations. The inner linear system solve was terminated when a relative residual tolerance of $\frac{\|r_{k}\|}{\|r_{0}\|}<10^{-5}$ was reached, where the norms used were the ones corresponding to results reported in the table: $\ell_{2}$-norm on the left and $H$-norm on the right.
For the diagonal preconditioner, we have observed experimentally for the smaller-size problems that the parameters of Lemma 2.4 satisfy $b\approx 9.18$, $c\approx 0.08$, and $bc<1$ in the $H$-norm. For the upper triangular preconditioner, $b\approx 8.28$, $c\approx 0.11$ and $bc<1$.
@@ 5 Concluding Remarks
When applying Theorem 1.5, a disk must be excluded from the field of values, and the remaining part should not surround the origin, as we have illustrated in Figure 1. To accomplish this, we require the imaginary part of the FOV to be small enough, which means that for the nonsymmetric saddle-point systems we consider, the skew-symmetric part of the preconditioned operator needs to be small in norm.
A finer geometric study of the field of values, beyond bounding it just by using the imaginary axis, may allow for broadening the scope of problems for which our analysis is applicable, including additional types nonsymmetric saddle-point linear systems.
Acknowledgments
We are very grateful to two knowledgable referees whose thorough and helpful reviews have greatly improved the quality of this paper.